OpenStudy (anonymous):

solve the equation by factoring and using zero product property. be sure to put the equation in stand form first. x^2-19x+80=-8

OpenStudy (anonymous):

wow i cant get the -19

OpenStudy (anonymous):

mmmm so u cant translate the -8 to the right side and get x^2 -19x +88=0 is that right?

OpenStudy (anonymous):

yeah i think so

OpenStudy (aripotta):

x^2 - 19x + 88 = 0 (x - 8)(x - 11)